Find the slope and y-intercept of 5x-2y=3.
-2y=-5x+3
y=(-5x=3)/-2
y=5/2x-3/2
The slope is 5/2 and the y-intercept is -3/2.
Did I do this problem correctly? Is this the right answer?
Thanks alot for your help in advance. :)
Elaine

Answer by mananth(16946)

Find the slope and y-intercept of 5x-2y=3.
-2y=-5x+3
y=(-5x+3)/-2
y=5/2x-3/2
The slope is 5/2 and the y-intercept is -3/2.
You are right